OpenStudy (anonymous):

Use the formula 5x + 3y = 40 to find y if x is 2

OpenStudy (anonymous):

5 ( 2) + 3y = 40 10 + 3y = 40 3y = 40 - 10 3y = 30 y = 10

pooja195 (pooja195):

Correct

OpenStudy (crabbyoldgamer):

yes. Good work.